OpenStudy (anonymous):

Is 27.5 a rational number

jimthompson5910 (jim_thompson5910):

Are you able to represent 27.5 as a fraction of two whole numbers (where the denominator isn't 0) ?

jimthompson5910 (jim_thompson5910):

For instance, 0.2 is a rational number because 0.2 = 2/10 = 1/5